For the following equation: 3x – 2y = 11:
Find x if y = 2
Find y if x = 1/3
Present the solutions as ordered pairs.

Answer by checkley77(12844)

3x – 2y = 11
WHEN Y=2 THEN:
3X-2*2=11
3x-4=11
3X=11+4
3X=15
X=15/3
X=5 ans.
-----------------------
WHEN X=1/3
3*1/3-2Y=11
1-2Y=11
-2Y=11-1
-2Y=10
Y=10/-2
Y=-5 ans.